Summit Products, Inc. is interested in producing and selling an improved widget. Market research indicates that customers would be willing to pay $76 for such a widget and that 36,000 units could be sold each year at this price. The current cost to produce the widget is estimated to be $52.

1. If Summit Products requires a 25% return on sales to undertake production, what is the target cost for the new widget?

a. $57.00

b. $19.00

c. $63.00

d. $51.00

2. Summit has learned that a competitor plans to introduce a similar widget at a price of $66. In response, Summit may reduce its selling price to $66. If Summit requires a 25% return on sales, what is the target cost for the new widget?

a. $49.50

b. $16.50

c. $19.00

d. $66.00

3. At a price of $66, Summit's market research indicates that it can sell 46,000 units per year. Assuming Summit can reach its new target cost, how will Summit's profit at the $66 price compare to what it would have earned in the absence of the competitor's product?

a. Profit will be $75,000 higher

b. Profit will be $75,000 lower

c. Profit will be unaffected if Summit can reach the revised target cost

d. Profit will be $46,000 higher
Answer #1

1. If Summit Products requires a 25% return on sales to undertake production, what is the target cost for the new widget?

Selling Price = $76

Return on sale = $76 * 25% = $19

Target Cost = $76 - $19 = $57

Option a is correct

2. Summit has learned that a competitor plans to introduce a similar widget at a price of $66. In response, Summit may reduce its selling price to $66. If Summit requires a 25% return on sales, what is the target cost for the new widget?

Selling Price = $66

Return on sale = $66 * 25% = $16.5

Target Cost = $66 - $16.5 = $49.5

Option a is correct

3. At a price of $66, Summit's market research indicates that it can sell 46,000 units per year. Assuming Summit can reach its new target cost, how will Summit's profit at the $66 price compare to what it would have earned in the absence of the competitor's product?

(46000 ´[$66 - $49.5]) - (36000 ´[$76 - $57]) = $75000

a. Profit will be $75,000 higher